OpenStudy (anonymous):

which of the following ordered pairs is not a solution to the inequality "y less than or equal to -3x - 5" ? (-4,0) (0,0) (-3,0) (-2,0)

OpenStudy (anonymous):

(x, y) Plug in the values of each of the points till you find one that makes the inequality true. \[y \le -3x - 5\] (-4, 0)\[0 \le -3(-4) - 5\]\[0 \le 12 - 5\]\[0 \le 7\]Is 0 less than or equal to 7? Is the inequality true?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Is 0 less than 7

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, so (-4, 0) is a possible solution to the problem. We're looking for the one that does NOT fit in, so lets keep checking. (0, 0) \[0 \le -3(0) - 5 \]\[0 \le 0 - 5 \]\[0 \le -5 \] Is 0 less than 5?

OpenStudy (anonymous):

*Is 0 less than -5?